Circular No. 7401

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1747.0–2853 S. Campana, Brera Observatory, Milan; and G. L. Israel and L. Stella,

Rome Observatory, on a behalf of a larger collaboration, report on a BeppoSAX observation of the transient source SAX J1747.0–2853 (IAUC 7377) on Mar. 16.71–17.48 UT: “The source was determined to lie at α = 17^{h}47^{m}03^{s}.1, δ = −28^{o}52′30′′ (equinox 2000.0; error radius 30′′ at 90-percent confidence level). During the BeppoSAX observation, a type-I burst was observed, testifying to the presence of a weakly magnetized neutron star. Based on the peak flux, we estimate a distance of < ∼ 8000 pc. The source was clearly detected up to 200 keV in the PDS instrument, showing a heavily absorbed (10^{23} cm^{−2}) spectrum modeled with a blackbody (kT = 1 keV) plus a cut-off (∼ 200 keV) power law (photon index γ = 2.6). The absorbed 0.1–200-keV flux amounts to 2 × 10^{−9} erg cm^{−2} s^{−1}.”

XTE J1118+480 C. Mauche, Lawrence Livermore National Laboratory; R. Hynes and

P. Charles, University of Southampton; and C. Haswell, Open University, on behalf of the EUVE project at the University of California at Berkeley, report: “EUVE obtained a 20 500-s exposure of XTE J1118+480 coincident with RXTE and Hubble Space Telescope (HST) observations during Apr. 8.10–8.71 UT. The mean count rate in the DS photometer was ∼ 1 count/s and the extreme-ultraviolet (EUV) spectrum is apparent in the SW spectrometer from the left edge of the detector at ∼ 7.0 nm to ∼ 12.0 nm. A crude estimate of the flux density at 9 nm is 4 × 10^{−13} erg cm^{−2} s^{−1} Å^{−1}. This is the first EUV spectrum of a blackhole-candidate x-ray binary. Additional EUVE observations are planned during observations by RXTE on Apr. 13 (20 000 s) and by RXTE, Chandra, and the HST on Apr. 18 (100 000 s).”

SUPERNOVA 1999em IN NGC 1637 CCD magnitudes, unfiltered unless otherwise noted: 1999 Nov. 4.746

UT, 14.1 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan); 13.591, 14.0 (Yoshida and Kadota); 2000 Jan. 5.267, 14.50 (M. Martignoni, Milan, Italy); 5.756, V = 13.96 ± 0.03 (D. Hanžl, Brno, Czech Rep.); 15.765, 14.59 (Martignoni); 21.774, 14.56 (Martignoni); 30.408, 14.5 (Yoshida and Kadota); Feb. 5.471, 14.2 (Yoshida and Kadota); 12.440, 14.6 (Yoshida and Kadota); Mar. 13.75, 15.7 ± 0.1 (K. Sárneczky, K. Sziládi, and L. Kiss, Piszkéstető station, Konkoly Observatory).

2000 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green